# Predicted versatile topological nodal magnons in the icosahedral quasicrystal 1/1 approximants

### v2

@@ IV Summary
In summary, we have theoretically clarified possible manifestation of topological magnon excitations in the whirling-antiwhirling magnetic state in the icosahedral quasicrystal 1/1 ACs. On the basis of the LSW+MTQC analysis, we have found two distinct types of topological nodal magnons referred to as type-A and type-B. It has been revealed that the former originates purely from magnetocrystalline symmetry, while the latter requires an accidenal band inversion for their emergence. We have also clarified that the long-wavelength effective theory accurately captures the two distinct types of topological nodal magnons, which indicates that long-wavelength measurements are promising for investigating a variety of nodal line physics in ACs.
Finally, we note that the exploration of materials for QCs and ACs with magnetism is rapidly advanced. Even merely with combinations of elements, a vast number of possibilities can be explored. Moreover, their composition ratios can be continuously varied for each. Therefore, it can be said that the number of achievable compounds is immense. In this sense, this work, which reveals the possible emergence of topological magnons in ACs for the first time, paves the new pathway to researches in the field of magnetism in QCs and related materials.

### v3

@@ IV Summary
In summary, we have theoretically clarified possible manifestation of topological magnon excitations in the whirling-antiwhirling magnetic state in the icosahedral quasicrystal 1/1 ACs. On the basis of the LSW+MTQC analysis, we have found two distinct types of topological nodal magnons referred to as type-A and type-B. It has been revealed that the former originates purely from magnetocrystalline symmetry, while the latter requires an accidental band inversion for their emergence. We have also clarified that the long-wavelength effective theory accurately captures the two distinct types of topological nodal magnons, which indicates that long-wavelength measurements are promising for investigating a variety of nodal line physics in ACs.
Finally, we note that the exploration of materials for quasicrystal approximants with magnetism is rapidly advanced. Even merely with combinations of elements, a vast number of possibilities can be explored. Moreover, their composition ratios can be continuously varied for each. Thus, it can be said that the number of achievable compounds is literally diverse. In this sense, this work, which reveals the possible emergence of topological magnons in ACs for the first time, paves the new pathway to research in the field of magnetism in ACs.